### v2

# Quantum Teleportation Coexisting with Conventional Classical Communications in Optical Fiber

Abstract
The ability for quantum and classical networks to operate in the same optical fibers would aid the deployment of quantum network technology. However, quantum performance can be susceptible to noise photons generated by spontaneous Raman scattering of high-power coexisting classical light. Quantum teleportation is a fundamental operation in quantum networking, but has yet to be demonstrated in fibers populated with high data rate conventional optical signals. In this paper, we demonstrate a three-node quantum state teleportation system coexisting with 400-Gbps C-band classical communications in 30.2 km of fiber. To protect quantum fidelity, Raman noise rates are suppressed using optimized O-band quantum channels and filtering in multiple degrees of freedom. Fidelity is shown to be well maintained with elevated classical powers as high as 18.7 dBm, which could support multiple classical channels with many terabits/s aggregate data rates. These results show the feasibility of advanced quantum and classical network applications operating within a unified fiber infrastructure.
journal: opticajournal
1 Introduction
Figure 1: (a) Conceptual diagram of the experiment. Alice prepares a qubit $|\psi\rangle_{A}$ that she wishes to transfer onto one photon of an entangled Bell state photon pair at Bob via quantum teleportation. Alice encodes $|\psi\rangle_{A}$ onto a single photon, which is multiplexed into a fiber of length $L_{AC}$ to co-propagate with a high-rate classical communications signal to a node at Charlie. The classical signal by-passes Charlie’s node to continue propagating over a fiber of length $L_{BC}$. Two photons are prepared in an entangled Bell state at Bob’s node. One photon counter-propagates with the classical signal over this fiber to Charlie to undergo a Bell state measurement (BSM) with Alice’s qubit. After a BSM detection, Alice’s state is destroyed whilst the other entangled photon at Bob’s node is projected onto the state $\sigma|\psi\rangle_{A}$, where $\sigma$ is a unitary transformation unique to the BSM result. (b) Experimental implementation. Quantum and classical signals coexist in 30.2  km of fiber ($L_{AC}=15.2\text{ km}$, $L_{BC}=15.0\text{ km}$). Alice and Bob both generate photon pairs with wavelengths of 1290 nm and 1310 nm via non-degenerate spontaneous parametric down-conversion. Alice’s source heralds a single photon, whereas Bob prepares a two-photon entangled state. Qubits are encoded in polarization and the 1290-nm photons from each source are transmitted to Charlie for the BSM. The classical signal has a rate of 400-Gbps and is allocated to the 1547.32 nm C-band wavelength channel. This signal travels an additional 24-km of deployed fiber before and after multiplexing, giving a total classical link distance of 78.2 km (FPC = fiber polarization controller, DWDM = dense wavelength division multiplexer, FWDM = O-band/C-band WDM, PPLN = periodically poled lithium niobate waveguide, $\lambda/2$ = half-wave plate, $\lambda/4$ = quarter-wave plate, LP = linear polarizer, LCR = liquid crystal retarder, PBS = polarizing beam splitter, FBS = 50:50 fiber coupler, VDL = variable optical delay line, FBG = fiber Bragg grating, CIR = circulator, $\text{D}_{j}$ = superconducting nanowire single photon detector, C = common port, P = pass port, R = reflect port, EDFA = erbium doped fiber amplifier).
The optical fiber infrastructure and telecommunications technology that underlie the Internet have been widely adopted by researchers seeking to develop quantum networks capable of applications such as quantum-enhanced cryptography, sensing, and networked quantum computing [1, 2]. However, since the majority of the existing fiber infrastructure is populated with classical network signals, whether or not quantum networking can be realized on a large scale will depend on the ability for both quantum and classical networks to operate in the same fibers due to limited existing fiber connections and the high cost of new installation.
Quantum and classical signals can readily share a single fiber via standard wavelength division multiplexing. However, inelastic scattering of high-power classical light can generate noise photons that obscure the detection of typically sub-photon-level quantum signals. Spontaneous Raman scattering (SpRS) is most prominent due to its broadband spectrum that spans the telecom bands. Without careful design, classical signals require significant attenuation to perform quantum operations, resulting in a trade-off between achievable classical data rates or quantum fidelity.
Quantum-classical integration has been studied for various systems, such as weak coherent states (WCSs) [3, 4, 5, 6, 7, 8, 9, 10, 11, 12], entangled photon pairs [13, 14, 15, 16, 17, 18, 19, 20, 21, 22], Bell state measurements (BSMs) on WCSs [23, 24], continuous variables [25], and squeezed light [26]. However, many systems that are integral to future quantum network architectures, such as those utilizing quantum teleportation, remain unexplored. Using the non-local properties of entanglement, quantum teleportation enables the transfer of a quantum state between two distant physical systems without direct transmission of the systems themselves [27]. It plays a foundational role in advanced applications such as quantum relays [28], quantum repeaters [29], networking quantum computers [30], and other applications in quantum science and technology [31].
Recent years have seen impressive progress towards implementing teleportation-based systems in quantum-dedicated fiber networks [32, 33, 34, 35, 36, 37, 31], but teleportation over fibers carrying conventional classical communications signals has yet to be demonstrated. The ability to perform teleportation alongside modern classical networking without sacrificing fidelity would be an important step in ensuring advanced quantum networks can make full use of the existing fiber infrastructure.
In this paper, we demonstrate a three-node quantum state teleportation system operating over 30.2 km of optical fiber that simultaneously carries high-power C-band classical communications at a rate of 400 Gbps. Teleportation is performed via a joint BSM on a single photon and one member of an entangled Bell state photon pair [38]. Utilizing wavelength multiplexing of quantum and classical signals, a heralded single photon carrying a polarization encoded qubit and one photon of an entangled Bell state photon pair are distributed over roughly 15 km of fiber to undergo a BSM near the midpoint of the fiber, whilst classical light traverses the full 30.2 km link plus an additional 48 km of deployed fiber.
We implement multiple techniques to minimize the degradation of quantum fidelity by SpRS noise. At the quantum source nodes, we generate and distribute photons with O-band wavelengths <1300 nm. Due to the multi-mode SpRS spectrum, these are the telecom wavelengths with nearly the minimum SpRS photon generation probability from the C-band classical source [17]. At quantum receiver nodes, we employ narrow-band spectral filtering and coincidence detection temporal filtering to reject uncorrelated in-band noise photons. This combines various approaches used in high-power quantum-classical coexistence studies of WCSs [9, 10, 11, 12], entangled photons [17], and BSMs on two WCSs [24], but has not been investigated in the context of quantum teleportation.
We first demonstrate that our system’s exceptionally low noise rates allow for no significant impact on entanglement distribution to the BSM nor Hong-Ou-Mandel (HOM) interference [39] between independent photons, which are the underlying operations of the photonic quantum teleportation protocol [38]. We then perform quantum state teleportation, where we measure an average state transfer fidelity of $89.8\pm 3.1$% with elevated classical C-band powers as high as 74 mW (18.7 dBm) launched into the fiber link. We discuss methods to improve both raw teleportation fidelity and noise-resilience as well as the implications of our results for future quantum networking.
This experiment demonstrates approaches for integrating many key classical and quantum network components, including high-rate C-band classical signals and optical amplifiers, single photon and entangled photon pair distribution, and multi-qubit operations such as BSMs all operating simultaneously in the same fibers.
2 Experimental Design
A conceptual diagram of the experiment is shown in Fig. 1(a). Alice generates a single photon to encode the quantum state $|\psi\rangle_{A}=\alpha|0\rangle+\beta|1\rangle$ that she wishes to teleport to Bob. Alice’s photon is multiplexed into an optical fiber of length $L_{AC}$ to co-propagate with a classical communications signal to a BSM node at Charlie. The classical signal is de-multiplexed just prior to the BSM and then re-multiplexed to bypass Charlie’s node. The classical signal then travels along another fiber of length $L_{BC}$, counter-propagating with respect to one photon from an entangled Bell state photon pair generated at Bob’s node. During the BSM at Charlie, both photons are irreversibly destroyed by the detection whilst Bob’s other photon of the entangled pair is projected onto the state $|\psi\rangle_{B}=\sigma|\psi\rangle_{A}$, where $\sigma$ is a unitary operation that is unique to the BSM result and can be classically communicated to Bob to reconstruct Alice’s state [27].
Fig. 1(b) shows the physical realization of the experiment. In our case, Alice’s state is encoded onto a heralded single photon’s polarization, $L_{AC}$ is 15.2 km, and $L_{BC}$ is 15.0 km. Thus, the quantum state transfer distance, and the distance in which quantum and classical signals coexist, is 30.2 km.
The classical source consists of a 400-Gbps C-band transceiver operating at 1547.32 nm (Ciena WaveLogic 5 Nano 400ZR). To reduce the amount of SpRS noise photons that are generated into the quantum channels, we strategically place our quantum signals in the O-band so that they have an anti-Stokes frequency detuning from the C-band as well as suppressed Raman gain at the far offset detuning between the C- and O-bands [4]. Instead of the commonly used 1310-nm wavelength allocation [9, 10, 11, 12, 24], our transmitted quantum signals are at a wavelength of 1290 nm. This channel has roughly an order of magnitude lower SpRS noise due to the non-uniform SpRS spectrum across the O-band [17], resulting in a higher tolerance to C-band power.
Alice and Bob generate their respective quantum signals via type-0 cascaded second harmonic generation-spontaneous parametric down conversion (c-SHG-SPDC) photon pair generation [40]. Alice and Bob both have independent periodically-poled lithium niobate (PPLN) waveguides that are each phase matched for SHG at 650 nm, which pumps the SPDC process to generate a two-mode squeezed vacuum state with a non-degenerate joint spectrum centered around 1300 nm. This state approximates a single photon pair given the probability of generating a pair per mode is $\ll 1$. The cascaded second order nonlinear process is analogous to four-wave mixing, which we pump using a 1300-nm distributed feedback laser that is intensity modulated using a lithium niobate modulator to generate pulses with an approximate 65-ps temporal full-width at half maximum (FWHM) and a repetition rate of 500 MHz. This pump is then split and directed to Alice and Bob’s nodes.
Alice’s qubit is initially encoded onto a heralded single photon. Her PPLN waveguide is pumped directly to generate photon pairs, where the detection of one photon indicates the presence of a single photon in the other mode. After the waveguide, we use a 100-GHz 1300-nm dense wavelength division multiplexer (DWDM) as a notch filter for the pump, followed by a 1310-nm DWDM to separate each photon into different fibers via the DWDM’s pass and reflect ports. The 1310-nm photon is filtered by a 10.5-GHz FWHM bandwidth fiber Bragg grating (FBG) filter and subsequently detected by a single photon detector. Due to energy conservation, this heralds a single photon with a center wavelength of 1290 nm. Alice’s qubit is then encoded onto the horizontal (H) and vertical (V) components of the heralded photon’s polarization $|\psi\rangle_{A}=\alpha|H\rangle+\beta|V\rangle$ using polarization waveplates. This photon is then multiplexed using a standard O-band/C-band WDM to co-propagate with the 400-Gbps C-band classical signal over 15.2 km of spooled optical fiber (SMF-28(R) ULL) to Charlie.
Bob’s entangled photon pair source is designed similarly, except his PPLN waveguide is placed inside a polarization Sagnac loop to generate the two-photon quantum state $\frac{1}{\sqrt{2}}(|{H_{s}H_{i}}\rangle+e^{i\delta}|{V_{s}V_{i}}\rangle)$. Bob’s 1290-nm photon is transmitted to counter-propagate with the 400-Gbps signal over 15.0 km of spooled fiber (SMF-28(R) ULL) to Charlie. However, Bob’s 1310-nm photon is kept locally to act as the target photon for transferring Alice’s state to Bob’s node. This photon is sent to a free-space polarization module which first compensates rotations in the fibers and prepares the $|\Psi^{-}\rangle_{B}$ Bell state using alignment signals and coincidence detection [41], and then uses a quarter-waveplate, half-waveplate, and polarizing beam splitter (PBS) combination to allow for arbitrary projective measurements of the polarization of Bob’s target photon. Subsequently, this photon is filtered by a 9.8-GHz FBG filter and then detected by a single-photon detector.
Figure 2: (a) Noise photon single photon count rate in the Bell state measurement detectors $D_{1}$ and $D_{2}$ as a function of the 400-Gbps C-band signal’s launch power $P_{\text{cl}}$ injected into the 30.2-km fiber link. (b) Visibility of non-local two-photon interference fringes for Bob’s entanglement distribution over 15.0 km to the Bell state measurement node as a function of $P_{\text{cl}}$ (top) and the measured two-fold coincidence count (CC) interference fringes in the vertical (V) and anti-diagonal (A) bases for $P_{\text{cl}}$ = 74 mW (bottom). (c) Hong-Ou-Mandel interference visibility between heralded photons from Alice and Bob as a function of $P_{\text{cl}}$ (top) and an interference dip measurement when $P_{\text{cl}}$ = 74 mW (bottom).
At Charlie’s node, a BSM is performed on Alice and Bob’s incident polarization qubits using conditional HOM interference at a 50:50 beam splitter [38]. Since HOM interference only occurs between identical photons, the fidelity of the BSM is governed by their spatial and spectro-temporal indistinguishability. To achieve this, a single-mode fiber 50:50 coupler is used for spatial indistinguishability and a variable optical delay in Alice’s pump path is used to ensure that both photons arrive simultaneously at the coupler. Furthermore, the 1290-nm photons are filtered by 10.8-GHz FBGs. Together with the 1310-nm FBGs, this extends the photons’ coherence times relative to the pump’s and increases the probability that each photon is in a single identical spectral mode [42].
These narrow-band spectral filters also play a critical role in improving tolerance to classical power due to the SpRS noise photons being spectrally independent of the quantum signal. Another potential source of noise is residual C-band photons due to insufficient isolation when filtering the O-band quantum signals. This is prevented by placing two cascaded O-band/C-band WDMs before the FBGs and a 1290-nm DWDM, which altogether provides >190 dB isolation from C-band light.
To perform the BSM, we place two PBSs in each output arm of 50:50 coupler before the single photon detectors $D_{1}$ and $D_{2}$, which are set to project onto the orthogonal polarization states $|H\rangle_{D_{1}}$ and $|V\rangle_{D_{2}}$. A two-fold coincidence detection at opposite ports of the 50:50 splitter indicates that Alice and Bob’s photons were in the $|\Psi^{-}\rangle_{AB}$ Bell state due to HOM interference that occurs if they were in any of the three other Bell states [38]. Due to the non-local correlations of Bob’s entangled state $|\Psi^{-}\rangle_{B}$, a successful BSM detection results in the target photon at Bob’s node being equivalent to Alice’s input state $|\psi\rangle_{A}$ [38]. We note that teleportation via detecting $|\Psi^{-}\rangle_{AB}$ can be performed without the PBSs. However, basis-selective Bell state analyzers are most common as they can double the success probability of a BSM by also measuring the $|\Psi^{+}\rangle$ Bell state, which could be done here if detectors were added to the other output ports of the PBSs [43]. Thus, this design allows us to evaluate how coexisting classical signals impact BSM nodes most likely to be implemented in future networks. We also note the PBSs at the BSM naturally filter SpRS noise in the polarization degree of freedom. Due to polarization mode dispersion, SpRS photons generated over long fibers are largely unpolarized [44, 17] which means polarizing elements can reduce noise rates by roughly a factor of two in each detector.
[…]
Since the modulation of the classical signal is uncorrelated with the pulsed quantum system, SpRS photons are randomly distributed in time relative to the quantum signals. As a result, narrow detection windows around the expected quantum signal arrival times act as a temporal filter on the noise. Furthermore, the inherent temporal correlations in photon pair detection can provide another layer of timing information that can increase the signal-to-noise ratio compared to sources such as WCSs [45].
The quantum teleportation-classical coexistence study is conducted in a laboratory located at Northwestern University’s Evanston campus. However, the classical transceiver is located in a building at the Northwestern University Chicago campus, which requires transmission over a 24-km deployed fiber pair to and from the Evanston location to operate. Thus, the classical signal travels an additional 48 km of fiber giving a total link distance of 78.2 km. The transmission loss over the 30.2-km fiber is 4.9 dB at the classical signal’s 1547.32-nm wavelength and the 1290-nm quantum signals have a total loss of 10.1 dB (5.1 dB and 5.0 dB for Alice and Bob’s photons, respectively). At Charlie, the bypassing of the BSM by the classical signal using WDMs adds an additional 1.2 dB insertion loss.
The total loss when transmitting the classical signal over the full 78.2-km fiber is 22.8 dB. Due to a receiver sensitivity of -18 dBm, the transceiver’s output power of -9 dBm requires amplification for the system to operate error-free. To achieve this, we place two cascaded erbium doped fiber amplifiers (EDFA) at the Evanston location just before multiplexing into the 30.2-km link to co-propagate with Alice’s photon. The minimum power launched into the 30.2-km fiber to achieve error-free communications could be as low as $-3$ dBm (0.5 mW) for the receiver located 30.2 km + 24 km away from the EDFA, and -10.7 dBm (85 $\mu$W) if the receiver were instead located directly at the end of the 30.2-km link. However, in this study we explore elevated launch powers to emulate more hostile environments that could be encountered in the field such as those found in high-capacity long-distance DWDM classical network systems.
3 Experimental Results
3.1 Noise rates, entanglement distribution, and Hong-Ou-Mandel interference
We now evaluate the performance of our quantum system. First, we characterize how much noise is introduced into the detectors $D_{1}$ and $D_{2}$ as a function of the classical power $P_{\text{cl}}$ that is multiplexed into the 30.2-km fiber. Fig. 2(a) shows the single photon count rates as we increase $P_{\text{cl}}$, where the SNSPD dark count rates have been subtracted. The extra noise scales linearly with power as 79.0 counts/s/mW and 97.9 counts/s/mW in $D_{1}$ and $D_{2}$, respectively, which differ due to slightly different detection system efficiencies. These results demonstrate that SpRS noise is significantly suppressed, as even mW power levels introduce extraneous noise on the order of the SNSPD dark count level. Since the 400-Gbps system can operate at $P_{\text{cl}}>0.5$ mW, this means that the SNSPDs hardly notice any rise in background counts when the system is operating at the minimum necessary classical power.
As teleportation requires quantum entanglement as a resource and the ability to have indistinguishable photons for the BSM, we independently characterize the quality of Bob’s entanglement distribution to Charlie and HOM interference as we vary $P_{\text{cl}}$. We set Alice and Bob’s mean photon pairs per pulse and polarization qubit $\mu$ generated inside each PPLN waveguide to approximately $\mu_{A}$ = $0.018$ and $\mu_{B}$ = $0.013$, respectively, which are chosen to balance rates with reduced performance due to multi-pair emission during SPDC [46].
Fig. 2(b) shows how Bob’s entanglement source’s non-local two-photon interference fringe visibility in the vertical and anti-diagonal bases is impacted after distributing one photon over 15.0 km to Charlie as we increase the classical power launched into the full 30.2-km link. The visibility is defined as $V_{\rm ent}=(R_{\rm max}-R_{\rm min})/(R_{\rm max}+R_{\rm min})$, where $R_{\rm max(\rm min)}$ are the maximum (minimum) two-fold coincidence counts of the interference fringe. Fig. 2(b) shows no noticeable degradation of $V_{\rm ent}$ up to $P_{\text{cl}}=$ 74 mW, in which we measure visibilities of $V_{V}=97.5\pm 0.1\%$ and $V_{A}=95.3\pm 0.2\%$. All error bars in this paper are determined via the Monte-Carlo method assuming Poisson photon counting statistics. These results indicate that entanglement distribution is well preserved at high classical powers, and that visibility is mainly limited from unity due to multi-pair emission during SPDC. These values are also well above the $1/\sqrt{2}$ bound for demonstrating the non-local nature of quantum entanglement [47]. Thus, this state is more than sufficient to use as a resource for quantum teleportation.
Next, we characterize HOM interference between heralded photons from Alice and Bob’s sources as we increase $P_{\text{cl}}$. To ensure polarization indistinguishably, we align the PBSs in the BSM setup to instead project onto $|H\rangle_{D_{1}}|H\rangle_{D_{2}}$ and prepare Alice and Bob’s heralded photons in the state $|H\rangle$. We then record four-fold coincidences as we vary the relative time-of-arrival of each photon using Alice’s variable optical delay. The quality of the interference is determined by the visibility of the fringe $V_{\rm HOM}=(R_{\rm max}-R_{\rm min})/R_{\rm max}$, where $R_{\rm max(min)}$ are the maximum (minimum) four-fold coincidence rates of the HOM interference fringe. The results are shown in Fig. 2(c). We measure a visibility of $V_{\rm HOM}=(R_{\rm max}-R_{\rm min})/R_{\rm max}=82.9\pm 4.5\%$ without the classical source, which is comparable to the $V_{\rm HOM}=83.1\pm 1.4\%$ visibility we obtain without distribution over the 30.2-km fiber. We measure $V_{\rm HOM}=80.3\pm 3.8\%$ with $P_{\text{cl}}$ set to 74 mW, showing interference is not significantly degraded by high-power coexisting C-band light. Each visibility is well above the classical bound of 50% [48], indicating our photons are sufficiently indistinguishable for performing quantum teleportation.
Figure 3: Four-fold coincidence counts as we scan Bob’s polarization analyzer setting along the Bloch sphere for Alice transmitting $|{D}\rangle$ (blue) and $|A\rangle$ (red) when the launch power of the 400-Gbps classical signal into the 30.2-km fiber is set to 74 mW.
Figure 4: Real and imaginary components of the density matrices obtained by single-qubit quantum state tomography on Bob’s target photon, conditioned on a successful BSM, whilst the launch power of the 400-Gbps C-band signal into the 30.2-km fiber is set to 74 mW. (a-d) show the results for Alice transmitting $|\psi\rangle_{A}=|H\rangle$, $|V\rangle$, $|D\rangle$, and $|A\rangle$, respectively.
3.2 Quantum teleportation coexisting with high-rate classical communications
Looking at the performance of entanglement distribution and HOM interference, we find that the key operations underlying quantum teleportation are well maintained at high C-band power levels. To demonstrate that this translates to the ability to perform noise-resilient teleportation, we evaluate the teleportation of various qubits while simultaneously transmitting 74 mW of C-band classical power.
Fig. 3 shows four-fold coincidence counts as we scan Bob’s polarization analyzer setting along the Bloch sphere for Alice transmitting $|{D}\rangle=\frac{1}{\sqrt{2}}(|{H}\rangle+|{V}\rangle)$ and $|{A}\rangle=\frac{1}{\sqrt{2}}(|{H}\rangle-|{V}\rangle)$. The coincidences exhibit sinusoidal fringes that coincide with Alice’s prepared qubit. We obtain fringe visibilities of $V_{D}=81.3\pm 5.4\%$ and $V_{A}=74.7\pm 4.7\%$, both values being above the $V>1/3$ classical limit [49].
Next, we perform maximum likelihood quantum state tomography [50] on Bob’s target photon to reconstruct the density matrix $\rho_{B}$ of his quantum state conditioned on a BSM detection. Fig. 4 shows $\rho_{B}$ for Alice transmitting the states $|H\rangle$, $|V\rangle$, $|D\rangle$, and $|A\rangle$. The fidelity of $\rho_{B}$ to Alice’s ideal qubit $|\psi\rangle_{A}$ is determined by calculating $F_{\psi}=\langle\psi_{A}|\rho_{B}|\psi_{A}\rangle$. State transfer in the H/V basis does not require HOM interference due to the projection onto $|H\rangle_{D_{1}}|V\rangle_{D_{2}}$ in the Bell state analyzer [51], where raw fidelity is mainly limited from unity due to multi-photon pair degradation. In this basis, we measure $F_{H}=97.5\pm 1.2\%$ and $F_{V}=95.8\pm 2.5\%$ for $|{\psi}\rangle_{A}=|{H}\rangle$ and $|{V}\rangle$, respectively. However, transferring coherent superpositions of $|H\rangle$ and $|V\rangle$ requires interference and thus exhibit more decoherence due to an imperfect spectro-temporal indistinguishability between Alice and Bob’s photons. For Alice transmitting $|{\psi}\rangle_{A}=|{D}\rangle$ and $|{A}\rangle$, we obtain fidelities of $F_{D}=87.5\pm 3.9\%$ and $F_{A}=85.5\pm 3.7\%$, respectively. Knowing that SpRS noise rates are independent of Alice’s prepared qubit, assuming symmetry of the equatorial states we calculate the average fidelity for an arbitrary qubit via $F_{\rm avg}=(F_{H}+F_{V}+2(F_{D}+F_{A}))/6$ in which we obtain $F_{\rm avg}=89.9\pm 3.1\%$. Furthermore, we measure $F_{\rm avg}=90.4\pm 2.7\%$ without sending the photons over the 30.2-km link, demonstrating that there is minimal degradation of fidelity from distribution over the fiber link nor the inclusion of the classical light. All of the measured fidelities are shown in Fig. 5. Since all are well above the classical bound of $F=2/3$ [49], we have clearly demonstrated that non-classical teleportation can be performed alongside high-rate conventional classical communications.
Figure 5: Fidelity of the reconstructed density matrix of Bob’s state to Alice’s initially prepared state for four different qubits as well as the average fidelity when the power of the 400-Gbps classical signal is set to 74 mW. The horizontal line shows the $F>2/3$ limit for demonstrating non-classical teleportation.
4 Discussion and Outlook
Since teleportation was conducted alongside much higher powers than necessary for the 400-Gbps classical system, this indicates that substantially higher data rates could have propagated through the fiber. The classical power used was 148 times greater than necessary for a single 400-Gbps channel, and 870 times greater than if the receiver were located directly at the end of the 30.2-km fiber. Thus, a multi-channel WDM classical system could easily have carried many terabits/s aggregate data rates without having a significant impact on quantum fidelity. The demonstrated powers are also on the order of the highest achieved in state-of-the-art single [10, 11, 12] and entangled [17] photon coexistence studies. This indicates the potential for teleportation-based quantum networking systems, such as quantum repeaters, operating alongside classical backbone networks [10].
Since our system uses two different photon pair sources, multiple nodes, and four-fold coincidence detection, our results indicate the potential for more complex systems operating alongside classical networks. For example, simple modifications to Alice’s source could allow entangled photon pair generation. A BSM at Charlie would then result in entanglement swapping [52], establishing entanglement between initially independent photons at Alice and Bob’s nodes. Due to our system’s high tolerance to classical power, it is reasonable to conclude that a similarly designed entanglement swapping system could operate alongside high-rate coexisting classical communications. Further, our multi-node configuration with the BSM performed near the midpoint of a longer fiber link is a key element for multi-node networking with a central quantum processor [34] and to implement quantum repeaters [29]. We note that these applications will also require quantum memories to store incident photons, the degree to which SpRS noise will impact various platforms and applications will require future investigation.
The state teleportation fidelity we obtained is comparable to recent experiments conducted over quantum-dedicated fiber networks [34, 33, 37], but could be further improved via standard methods. We showed that opposed to SpRS noise from coexisting classical light, our fidelity was mainly limited by multi-photon pair emission during SPDC and imperfect photon indistinguishably during the BSM. Raw fidelity could thus be improved by reducing the photon pair generation probability [46] and increasing indistinguishability by decreasing the ratio between the source’s pump spectral bandwidth to the photon filter bandwidths [42]. When considering SpRS noise, narrower filters would simultaneously increase indistinguishability and reject more SpRS noise. Altogether, employing these approaches could allow our teleportation fidelity to approach more closely to unity and have an even higher tolerance to classical power.
Finally, classical signals are often used for quantum networking purposes. Some applications include communicating measurement results between nodes (eg., the BSM), synchronizing distant pulsed sources and detectors to a common clock, and channel monitoring. Similarly, quantum-classical coexistence can enhance functionality without sacrificing potentially limited fiber resources [16, 53, 22, 21]. Since these signals can typically operate well below the powers demonstrated here (sub-mW levels), our results show that including these signals in our system would have negligible impact if allocated to the C- or L-bands.
In conclusion, we have demonstrated a three-node quantum state teleportation system operating in a 30.2-km fiber link that simultaneously carries 400-Gbps conventional classical communications. By employing various methods to suppress the impact of SpRS noise photons, teleportation fidelity was well maintained with elevated classical powers capable of transmitting many terabits/s aggregate data rates. Since we have clearly shown that appropriately designed teleportation systems can operate alongside modern high-rate data traffic, this work demonstrates a significant step towards ensuring future quantum networking can be realized anywhere in the existing fiber infrastructure.
[…]
Funding This work is funded by Subcontract No. 664603 from Fermi Research Alliance, LLC (FRA) to Northwestern University issued under Prime Contract No. DE-AC02-07CH11359 between FRA and the U.S. Department of Energy (DOE). Although the work is supported by the DOE’s Advanced Scientific Computing Research Transparent Optical Quantum Networks for Distributed Science program, no government endorsement is implied.
[…]
Acknowledgments The authors would like to acknowledge the support of StarLight International/National Communications Exchange Facility and Ciena Corporation. We further thank the members of the Advanced Quantum Networks for Scientific Discovery (AQNET-SD) project, which is a collaboration between Northwestern University, Caltech, Fermilab National Laboratory, Argonne National Laboratory, the Jet Propulsion Laboratory, and University of Illinois Urbana-Champaign towards advancing quantum network technology and applications.
[…]
Disclosures The authors declare no conflicts of interest.
[…]
Data Availability Statement Data may be obtained from the authors upon reasonable request.

### v3

@@
-Quantum Teleportation Coexisting with Conventional Classical Communications in Optical Fiber
-Abstract
-The ability for quantum and classical networks to operate in the same optical fibers would aid the deployment of quantum network technology.
-However, quantum performance can be susceptible to noise photons generated by spontaneous Raman scattering of high-power coexisting classical light.
-Quantum teleportation is a fundamental operation in quantum networking, but has yet to be demonstrated in fibers populated with high data rate conventional optical signals.
-In this paper, we demonstrate a three-node quantum state teleportation system coexisting with 400-Gbps C-band classical communications in 30.2 km of fiber.
-To protect quantum fidelity, Raman noise rates are suppressed using optimized O-band quantum channels and filtering in multiple degrees of freedom.
-Fidelity is shown to be well maintained with elevated classical powers as high as 18.7 dBm, which could support multiple classical channels with many terabits/s aggregate data rates.
+Quantum Teleportation Coexisting with Classical Communications in Optical Fiber
+Quantum Teleportation Coexisting with Classical Communications in Optical Fiber
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+journal: opticajournalarticletype: Research Article
+{abstract*}
+The ability for quantum and conventional networks to operate in the same optical fibers would aid the deployment of quantum network technology on a large scale.
+Quantum teleportation is a fundamental operation in quantum networking, but has yet to be demonstrated in fibers populated with high-power conventional optical signals.
+Here we report to the best of our knowledge the first demonstration of quantum teleportation over fibers carrying conventional telecommunications traffic.
+Quantum state transfer is achieved over a 30.2-km fiber carrying 400-Gbps C-band classical traffic with a Bell state measurement performed at the fiber’s midpoint.
+To protect quantum fidelity from spontaneous Raman scattering noise, we use optimal O-band quantum channels, narrow spectro-temporal filtering, and multi-photon coincidence detection.
+Fidelity is shown to be well maintained with an elevated C-band classical power of 18.7 dBm, which could support multiple classical channels totaling many terabits/s aggregate data rates.
 These results show the feasibility of advanced quantum and classical network applications operating within a unified fiber infrastructure.
-journal: opticajournal
 1 Introduction
-Figure 1: (a) Conceptual diagram of the experiment.
-Alice prepares a qubit $|\psi\rangle_{A}$ that she wishes to transfer onto one photon of an entangled Bell state photon pair at Bob via quantum teleportation.
-Alice encodes $|\psi\rangle_{A}$ onto a single photon, which is multiplexed into a fiber of length $L_{AC}$ to co-propagate with a high-rate classical communications signal to a node at Charlie.
-The classical signal by-passes Charlie’s node to continue propagating over a fiber of length $L_{BC}$.
-Two photons are prepared in an entangled Bell state at Bob’s node.
-One photon counter-propagates with the classical signal over this fiber to Charlie to undergo a Bell state measurement (BSM) with Alice’s qubit.
-After a BSM detection, Alice’s state is destroyed whilst the other entangled photon at Bob’s node is projected onto the state $\sigma|\psi\rangle_{A}$, where $\sigma$ is a unitary transformation unique to the BSM result.
-(b) Experimental implementation.
-Quantum and classical signals coexist in 30.2  km of fiber ($L_{AC}=15.2\text{ km}$, $L_{BC}=15.0\text{ km}$).
-Alice and Bob both generate photon pairs with wavelengths of 1290 nm and 1310 nm via non-degenerate spontaneous parametric down-conversion.
-Alice’s source heralds a single photon, whereas Bob prepares a two-photon entangled state.
-Qubits are encoded in polarization and the 1290-nm photons from each source are transmitted to Charlie for the BSM.
-The classical signal has a rate of 400-Gbps and is allocated to the 1547.32 nm C-band wavelength channel.
-This signal travels an additional 24-km of deployed fiber before and after multiplexing, giving a total classical link distance of 78.2 km (FPC = fiber polarization controller, DWDM = dense wavelength division multiplexer, FWDM = O-band/C-band WDM, PPLN = periodically poled lithium niobate waveguide, $\lambda/2$ = half-wave plate, $\lambda/4$ = quarter-wave plate, LP = linear polarizer, LCR = liquid crystal retarder, PBS = polarizing beam splitter, FBS = 50:50 fiber coupler, VDL = variable optical delay line, FBG = fiber Bragg grating, CIR = circulator, $\text{D}_{j}$ = superconducting nanowire single photon detector, C = common port, P = pass port, R = reflect port, EDFA = erbium doped fiber amplifier).
 The optical fiber infrastructure and telecommunications technology that underlie the Internet have been widely adopted by researchers seeking to develop quantum networks capable of applications such as quantum-enhanced cryptography, sensing, and networked quantum computing [1, 2].
-However, since the majority of the existing fiber infrastructure is populated with classical network signals, whether or not quantum networking can be realized on a large scale will depend on the ability for both quantum and classical networks to operate in the same fibers due to limited existing fiber connections and the high cost of new installation.
-Quantum and classical signals can readily share a single fiber via standard wavelength division multiplexing.
-However, inelastic scattering of high-power classical light can generate noise photons that obscure the detection of typically sub-photon-level quantum signals.
-Spontaneous Raman scattering (SpRS) is most prominent due to its broadband spectrum that spans the telecom bands.
-Without careful design, classical signals require significant attenuation to perform quantum operations, resulting in a trade-off between achievable classical data rates or quantum fidelity.
-Quantum-classical integration has been studied for various systems, such as weak coherent states (WCSs) [3, 4, 5, 6, 7, 8, 9, 10, 11, 12], entangled photon pairs [13, 14, 15, 16, 17, 18, 19, 20, 21, 22], Bell state measurements (BSMs) on WCSs [23, 24], continuous variables [25], and squeezed light [26].
-However, many systems that are integral to future quantum network architectures, such as those utilizing quantum teleportation, remain unexplored.
-Using the non-local properties of entanglement, quantum teleportation enables the transfer of a quantum state between two distant physical systems without direct transmission of the systems themselves [27].
+However, since the majority of the existing fiber infrastructure is populated with conventional telecommunications traffic and due to the economic cost of leasing or installing new fiber, whether quantum networking can be realized on a large scale will depend on the ability to propagate quantum signals in the same fiber as high-power classical signals.
+Quantum and classical signals can readily share a single fiber via wavelength division multiplexing.
+However, it was first shown in [3] that noise photons from inelastic scattering of the high-power classical light can obscure the detection of often sub-photon-level quantum signals.
+Of the potential sources of noise, spontaneous Raman scattering (SpRS) is most dominant due to its broadband spectrum meaning noise can never be entirely prevented [4].
+Without careful design, a trade-off between the capabilities of conventional and quantum networking arises.
+The study of quantum-classical coexistence has a long history with many experiments conducted for applications using weak coherent state (WCS) sources [3, 4, 5, 6, 7, 8, 9, 10, 11, 12], entangled photon pairs [13, 14, 15, 16, 17, 18, 19, 20, 21, 22], Bell state measurements (BSMs) on WCSs [23, 24], continuous variables [25], and squeezed light [26].
+Although significant achievements and insights have been made in these studies, all experiments to date have focused on systems that directly transmit quantum information between network nodes.
+However, many next-generation quantum applications require the disembodied transfer of quantum states between users.
+Enabled by the non-local properties of quantum entanglement, quantum teleportation allows the transfer of a quantum state between two distant physical systems without the need for direct transmission [27].
 It plays a foundational role in advanced applications such as quantum relays [28], quantum repeaters [29], networking quantum computers [30], and other applications in quantum science and technology [31].
-Recent years have seen impressive progress towards implementing teleportation-based systems in quantum-dedicated fiber networks [32, 33, 34, 35, 36, 37, 31], but teleportation over fibers carrying conventional classical communications signals has yet to be demonstrated.
-The ability to perform teleportation alongside modern classical networking without sacrificing fidelity would be an important step in ensuring advanced quantum networks can make full use of the existing fiber infrastructure.
+Recent years have seen impressive progress towards implementing teleportation-based networks in quantum-dedicated fiber [32, 33, 34, 35, 36, 37, 38, 39, 31], but teleportation over fibers carrying conventional communications has yet to be demonstrated.
+As a result, many open questions on the feasibility, limitations, and potential benefits of teleportation in quantum-classical networking remain unexplored.
 In this paper, we demonstrate a three-node quantum state teleportation system operating over 30.2 km of optical fiber that simultaneously carries high-power C-band classical communications at a rate of 400 Gbps.
@@
 Utilizing wavelength multiplexing of quantum and classical signals, a heralded single photon carrying a polarization encoded qubit and one photon of an entangled Bell state photon pair are distributed over roughly 15 km of fiber to undergo a BSM near the midpoint of the fiber, whilst classical light traverses the full 30.2 km link plus an additional 48 km of deployed fiber.
-We implement multiple techniques to minimize the degradation of quantum fidelity by SpRS noise.
-At the quantum source nodes, we generate and distribute photons with O-band wavelengths <1300 nm.
-Due to the multi-mode SpRS spectrum, these are the telecom wavelengths with nearly the minimum SpRS photon generation probability from the C-band classical source [17].
-At quantum receiver nodes, we employ narrow-band spectral filtering and coincidence detection temporal filtering to reject uncorrelated in-band noise photons.
-This combines various approaches used in high-power quantum-classical coexistence studies of WCSs [9, 10, 11, 12], entangled photons [17], and BSMs on two WCSs [24], but has not been investigated in the context of quantum teleportation.
-We first demonstrate that our system’s exceptionally low noise rates allow for no significant impact on entanglement distribution to the BSM nor Hong-Ou-Mandel (HOM) interference [39] between independent photons, which are the underlying operations of the photonic quantum teleportation protocol [38].
-We then perform quantum state teleportation, where we measure an average state transfer fidelity of $89.8\pm 3.1$% with elevated classical C-band powers as high as 74 mW (18.7 dBm) launched into the fiber link.
-We discuss methods to improve both raw teleportation fidelity and noise-resilience as well as the implications of our results for future quantum networking.
-This experiment demonstrates approaches for integrating many key classical and quantum network components, including high-rate C-band classical signals and optical amplifiers, single photon and entangled photon pair distribution, and multi-qubit operations such as BSMs all operating simultaneously in the same fibers.
+We implement multiple techniques to mitigate degradation due to noise photons.
+At the quantum source nodes, we make use of the low SpRS noise in the O-band from strong C-band light [3], which has been shown successful for protecting quantum states from high-power C-band communications [9, 10, 11, 12, 24, 17].
+More specifically, we allocate coexisting photons to the 1290-nm quantum channels to have nearly the lowest SpRS generation probability achievable using telecom-band wavelength engineering [17].
+At quantum receivers, we use narrow-band spectro-temporal filtering to reject the detection of uncorrelated SpRS photons.
+Furthermore, each source uses spontaneous parametric down-conversion to generate photon pairs with strong temporal correlations to further reject noise via coincidence detection of the four photons.
+The design is also similar to a quantum relay, which has been predicted to reduce the impact of detector dark count noise [28] indicating a potential application in quantum-classical coexistence.
+We first demonstrate that our approach allows high-power C-band communications to have no significant impact on entanglement distribution to the BSM node nor Hong-Ou-Mandel (HOM) interference [41] between independent heralded photons, which are the underlying operations of the quantum teleportation protocol [40].
+We then perform quantum state teleportation and demonstrate that fidelity is maintained at elevated C-band launch powers as high as 74 mW (18.7 dBm), which is on the order of the highest achieved using direct transmission of quantum states [10, 11, 12, 17, 25].
+Using a theoretical model, we then discuss the the pros and cons of using noisier quantum wavelength channels (eg., C-band), the effects of storage in imperfect quantum memories, and compare to methods using direct qubit transmission.
+This experiment demonstrates approaches for integrating many key classical and quantum network components, including high-rate classical signals, optical amplifiers, single and entangled photon pair distribution, and multi-qubit operations such as BSMs all operating simultaneously in the same fibers.
+Figure 1: (a) Conceptual diagram of the experiment.
+Alice prepares a qubit $|\psi\rangle_{A}$ that she wishes to transfer onto a photon at Bob’s node via quantum state teleportation.
+Alice encodes $|\psi\rangle_{A}$ onto a single photon, which is wavelength-division multiplexed into a fiber of length $L_{AC}$ to co-propagate with conventional classical communications traffic to a node at Charlie.
+The classical signal is de-multiplexed to bypass Charlie’s node and re-multiplexed to continue propagating over a fiber of length $L_{BC}$.
+Two photons generated at Bob’s node are prepared in an entangled Bell state, where one is transmitted to counter-propagate with the classical light to Charlie to undergo a Bell state measurement (BSM) with Alice’s qubit.
+After a BSM detection, Alice’s state is destroyed whilst the photon kept at Bob’s node is projected onto the state $\sigma|\psi\rangle_{A}$, where $\sigma$ is a unitary Pauli transformation unique to the BSM result.
+(b) Experimental implementation.
+Quantum and classical signals coexist in 30.2  km of fiber ($L_{AC}=15.2\text{ km}$, $L_{BC}=15.0\text{ km}$).
+Alice and Bob both generate photon pairs via non-degenerate spontaneous parametric down-conversion with center wavelengths of 1290 nm and 1310 nm.
+Alice’s source heralds a single photon, whereas Bob prepares a two-photon entangled Bell state.
+Qubits are encoded in polarization and the 1290-nm photons from each source are transmitted to Charlie for the BSM and teleportation is evaluated via four-fold coincidence detection and polarization analysis at Bob.
+The classical signal has a rate of 400-Gbps and is allocated to the 1547.32 nm C-band wavelength channel.
+This signal travels an additional 24-km of deployed fiber before and after multiplexing giving a total classical link distance of 78.2 km (FPC = fiber polarization controller, DWDM = dense wavelength division multiplexer, FWDM = O-band/C-band WDM, PPLN = periodically poled lithium niobate waveguide, $\lambda/2$ = half-wave plate, $\lambda/4$ = quarter-wave plate, LP = linear polarizer, LCR = liquid crystal retarder, PBS = polarizing beam splitter, FBS = 50:50 fiber coupler, VDL = variable optical delay line, FBG = fiber Bragg grating, CIR = circulator, $\text{D}_{j}$ = superconducting nanowire single photon detector, C = common port, P = pass port, R = reflect port, EDFA = erbium doped fiber amplifier).
 2 Experimental Design
@@
 The classical signal then travels along another fiber of length $L_{BC}$, counter-propagating with respect to one photon from an entangled Bell state photon pair generated at Bob’s node.
-During the BSM at Charlie, both photons are irreversibly destroyed by the detection whilst Bob’s other photon of the entangled pair is projected onto the state $|\psi\rangle_{B}=\sigma|\psi\rangle_{A}$, where $\sigma$ is a unitary operation that is unique to the BSM result and can be classically communicated to Bob to reconstruct Alice’s state [27].
+During the BSM at Charlie, both photons are irreversibly destroyed by the detection whilst Bob’s other photon of the entangled pair is projected onto the state $|\psi\rangle_{B}=\sigma|\psi\rangle_{A}$, where $\sigma$ is a unitary operation that is unique to the BSM result and can be classically communicated to Bob to recover Alice’s state [27] either by physically applying the unitary or accounting for it in data post-processing.
 Fig.
@@
 The classical source consists of a 400-Gbps C-band transceiver operating at 1547.32 nm (Ciena WaveLogic 5 Nano 400ZR).
-To reduce the amount of SpRS noise photons that are generated into the quantum channels, we strategically place our quantum signals in the O-band so that they have an anti-Stokes frequency detuning from the C-band as well as suppressed Raman gain at the far offset detuning between the C- and O-bands [4].
-Instead of the commonly used 1310-nm wavelength allocation [9, 10, 11, 12, 24], our transmitted quantum signals are at a wavelength of 1290 nm.
-This channel has roughly an order of magnitude lower SpRS noise due to the non-uniform SpRS spectrum across the O-band [17], resulting in a higher tolerance to C-band power.
-Alice and Bob generate their respective quantum signals via type-0 cascaded second harmonic generation-spontaneous parametric down conversion (c-SHG-SPDC) photon pair generation [40].
-Alice and Bob both have independent periodically-poled lithium niobate (PPLN) waveguides that are each phase matched for SHG at 650 nm, which pumps the SPDC process to generate a two-mode squeezed vacuum state with a non-degenerate joint spectrum centered around 1300 nm.
-This state approximates a single photon pair given the probability of generating a pair per mode is $\ll 1$.
-The cascaded second order nonlinear process is analogous to four-wave mixing, which we pump using a 1300-nm distributed feedback laser that is intensity modulated using a lithium niobate modulator to generate pulses with an approximate 65-ps temporal full-width at half maximum (FWHM) and a repetition rate of 500 MHz.
-This pump is then split and directed to Alice and Bob’s nodes.
+To reduce the amount of SpRS noise photons that are generated into the quantum channels, we generate photons at O-band wavelengths so that they have an anti-Stokes frequency detuning from the C-band light as well as suppressed Raman gain at the far offset detuning between the C- and O-bands [4].
+Instead of the most commonly used 1310-nm O-band channel [9, 10, 11, 12, 24], we suppress SpRS by roughly another order of magnitude using the 1290-nm channel due to the multi-mode SpRS noise spectrum from C-band light [17].
+Both Alice and Bob generate their respective quantum signals via type-0 cascaded second harmonic generation-spontaneous parametric down conversion (c-SHG-SPDC) photon pair generation [42].
+Alice and Bob each have independent periodically-poled lithium niobate (PPLN) waveguides that are phase matched for SHG at 650 nm.
+The cascaded second-order nonlinear process is analogous to four-wave mixing, which we pump using a 1300-nm continuous-wave distributed feedback laser that is intensity modulated using a lithium niobate Mach-Zehnder modulator to generate pulses with an approximate 65-ps temporal full-width at half maximum (FWHM) and a pulse repetition rate of 500 MHz.
+This pump is then split and directed to Alice and Bob’s nodes (see supp.
+materials).
+The 650-nm pulses generated inside each waveguide simultaneously drives SPDC to produce a two-mode squeezed vacuum state with a non-degenerate joint spectrum centered around 1300 nm.
+This state approximates a probabilistic source of single-photon pairs given the probability of generating a pair per pulse is $\ll 1$.
 Alice’s qubit is initially encoded onto a heralded single photon.
-Her PPLN waveguide is pumped directly to generate photon pairs, where the detection of one photon indicates the presence of a single photon in the other mode.
+Her PPLN waveguide is pumped directly to generate photon pairs.
 After the waveguide, we use a 100-GHz 1300-nm dense wavelength division multiplexer (DWDM) as a notch filter for the pump, followed by a 1310-nm DWDM to separate each photon into different fibers via the DWDM’s pass and reflect ports.
-The 1310-nm photon is filtered by a 10.5-GHz FWHM bandwidth fiber Bragg grating (FBG) filter and subsequently detected by a single photon detector.
-Due to energy conservation, this heralds a single photon with a center wavelength of 1290 nm.
+The 1310-nm photon is filtered by a fiber Bragg grating (FBG) with FWHM bandwidth of 60 pm and then detected by a superconducting nanowire single-photon detector (SNSPD).
+Due to energy conservation, this detection heralds the presence of a photon with a center wavelength of 1290 nm in the other mode.
 Alice’s qubit is then encoded onto the horizontal (H) and vertical (V) components of the heralded photon’s polarization $|\psi\rangle_{A}=\alpha|H\rangle+\beta|V\rangle$ using polarization waveplates.
-This photon is then multiplexed using a standard O-band/C-band WDM to co-propagate with the 400-Gbps C-band classical signal over 15.2 km of spooled optical fiber (SMF-28(R) ULL) to Charlie.
-Bob’s entangled photon pair source is designed similarly, except his PPLN waveguide is placed inside a polarization Sagnac loop to generate the two-photon quantum state $\frac{1}{\sqrt{2}}(|{H_{s}H_{i}}\rangle+e^{i\delta}|{V_{s}V_{i}}\rangle)$.
+This photon is then multiplexed using an O-band/C-band WDM to co-propagate with the 400-Gbps C-band classical signal over 15.2 km of spooled optical fiber (SMF-28(R) ULL) to Charlie.
+Bob’s source is designed similarly, except his PPLN waveguide is placed inside a polarization Sagnac loop to generate polarization entangled photon pairs using c-SHG-SPDC [42].
 Bob’s 1290-nm photon is transmitted to counter-propagate with the 400-Gbps signal over 15.0 km of spooled fiber (SMF-28(R) ULL) to Charlie.
-However, Bob’s 1310-nm photon is kept locally to act as the target photon for transferring Alice’s state to Bob’s node.
-This photon is sent to a free-space polarization module which first compensates rotations in the fibers and prepares the $|\Psi^{-}\rangle_{B}$ Bell state using alignment signals and coincidence detection [41], and then uses a quarter-waveplate, half-waveplate, and polarizing beam splitter (PBS) combination to allow for arbitrary projective measurements of the polarization of Bob’s target photon.
-Subsequently, this photon is filtered by a 9.8-GHz FBG filter and then detected by a single-photon detector.
-Figure 2: (a) Noise photon single photon count rate in the Bell state measurement detectors $D_{1}$ and $D_{2}$ as a function of the 400-Gbps C-band signal’s launch power $P_{\text{cl}}$ injected into the 30.2-km fiber link.
-(b) Visibility of non-local two-photon interference fringes for Bob’s entanglement distribution over 15.0 km to the Bell state measurement node as a function of $P_{\text{cl}}$ (top) and the measured two-fold coincidence count (CC) interference fringes in the vertical (V) and anti-diagonal (A) bases for $P_{\text{cl}}$ = 74 mW (bottom).
-(c) Hong-Ou-Mandel interference visibility between heralded photons from Alice and Bob as a function of $P_{\text{cl}}$ (top) and an interference dip measurement when $P_{\text{cl}}$ = 74 mW (bottom).
+However, Bob’s 1310-nm photon is kept locally to act as the target photon for teleporting Alice’s state to Bob’s node.
+This photon is sent to a free-space polarization module which compensates polarization rotations in the fibers and prepares Bob’s photons in the $|\Psi^{-}\rangle_{B}$ Bell state using alignment signals and coincidence detection [43], and then uses a quarter-waveplate, half-waveplate, and polarizing beam splitter (PBS) combination to characterize the polarization of Bob’s target photon.
+Subsequently, this photon is filtered by a 56 pm bandwidth FBG filter and then detected by an SNSPD.
 At Charlie’s node, a BSM is performed on Alice and Bob’s incident polarization qubits using conditional HOM interference at a 50:50 beam splitter [40].
@@
 To achieve this, a single-mode fiber 50:50 coupler is used for spatial indistinguishability and a variable optical delay in Alice’s pump path is used to ensure that both photons arrive simultaneously at the coupler.
-Furthermore, the 1290-nm photons are filtered by 10.8-GHz FBGs.
-Together with the 1310-nm FBGs, this extends the photons’ coherence times relative to the pump’s and increases the probability that each photon is in a single identical spectral mode [42].
-These narrow-band spectral filters also play a critical role in improving tolerance to classical power due to the SpRS noise photons being spectrally independent of the quantum signal.
-Another potential source of noise is residual C-band photons due to insufficient isolation when filtering the O-band quantum signals.
-This is prevented by placing two cascaded O-band/C-band WDMs before the FBGs and a 1290-nm DWDM, which altogether provides >190 dB isolation from C-band light.
-To perform the BSM, we place two PBSs in each output arm of 50:50 coupler before the single photon detectors $D_{1}$ and $D_{2}$, which are set to project onto the orthogonal polarization states $|H\rangle_{D_{1}}$ and $|V\rangle_{D_{2}}$.
-A two-fold coincidence detection at opposite ports of the 50:50 splitter indicates that Alice and Bob’s photons were in the $|\Psi^{-}\rangle_{AB}$ Bell state due to HOM interference that occurs if they were in any of the three other Bell states [38].
-Due to the non-local correlations of Bob’s entangled state $|\Psi^{-}\rangle_{B}$, a successful BSM detection results in the target photon at Bob’s node being equivalent to Alice’s input state $|\psi\rangle_{A}$ [38].
-We note that teleportation via detecting $|\Psi^{-}\rangle_{AB}$ can be performed without the PBSs.
-However, basis-selective Bell state analyzers are most common as they can double the success probability of a BSM by also measuring the $|\Psi^{+}\rangle$ Bell state, which could be done here if detectors were added to the other output ports of the PBSs [43].
-Thus, this design allows us to evaluate how coexisting classical signals impact BSM nodes most likely to be implemented in future networks.
-We also note the PBSs at the BSM naturally filter SpRS noise in the polarization degree of freedom.
-Due to polarization mode dispersion, SpRS photons generated over long fibers are largely unpolarized [44, 17] which means polarizing elements can reduce noise rates by roughly a factor of two in each detector.
+Furthermore, the 1290-nm photons are filtered by 60-pm bandwidth FBGs.
+Together with the 1310-nm FBGs, this extends the photons’ coherence times relative to the pump’s to increase the probability that each photon is in a single identical spectral mode [44].
+These narrow-band filters are also crucial for improving tolerance to classical power due to the spectral independence between the SpRS noise photons and the quantum signals.
+To prevent the detection of C-band photons due to insufficient filter isolation, we cascade two cascaded O-band/C-band WDMs before the FBGs and a 1290-nm DWDM to achieve >190 dB rejection of C-band light.
+To perform the BSM, we place PBSs in each output of the 50:50 coupler before the detectors $D_{1}$ and $D_{2}$ which are set to project onto orthogonal polarizations $|H\rangle_{D_{1}}$ and $|V\rangle_{D_{2}}$.
+A two-fold coincidence detection then indicates that Alice and Bob’s photons were in the $|\Psi^{-}\rangle_{AB}$ Bell state due to HOM interference effects on the other Bell states [40].
+Due to the non-local correlations of Bob’s entangled photons in the state $|\Psi^{-}\rangle_{B}$, the destructive measurement of $|\Psi^{-}\rangle_{AB}$ results in the remaining photon at Bob’s node being projected onto Alice’s initial prepared state $|\psi\rangle_{A}$ [40].
+Although teleportation can be performed without projecting onto the $H/V$ basis, Bell state analyzers that measure in the qubit’s eigenbasis can double the success probability of a BSM by also measuring the $|\Psi^{+}\rangle$ Bell state if detectors were added to the other ports of our PBSs [45].
+Including the PBSs thus allows us to investigate BSM nodes most likely to be implemented in future networks.
+We also note that due to polarization mode dispersion over long fibers, the SpRS noise is largely unpolarized [46, 17] which means the polarizing elements roughly halve the amount of SpRS entering each detector.
 […]
-Since the modulation of the classical signal is uncorrelated with the pulsed quantum system, SpRS photons are randomly distributed in time relative to the quantum signals.
-As a result, narrow detection windows around the expected quantum signal arrival times act as a temporal filter on the noise.
-Furthermore, the inherent temporal correlations in photon pair detection can provide another layer of timing information that can increase the signal-to-noise ratio compared to sources such as WCSs [45].
-The quantum teleportation-classical coexistence study is conducted in a laboratory located at Northwestern University’s Evanston campus.
-However, the classical transceiver is located in a building at the Northwestern University Chicago campus, which requires transmission over a 24-km deployed fiber pair to and from the Evanston location to operate.
-Thus, the classical signal travels an additional 48 km of fiber giving a total link distance of 78.2 km.
+Since the classical data modulation is uncorrelated with the pulsed quantum sources in the time-domain, the coincidence detection window acts as a temporal filter of the SpRS noise.
+The temporal correlations in photon pair generation can also be exploited for time-domain filtering.
+Heralding single photons can improve tolerance to noise compared to sources such as WCSs by accepting a detection only when the heralding signal ensures a photon had been generated which reduces the probability of transmitting a vacuum state [47].
+In our four-photon teleportation system, both Alice and Bob’s pair sources benefit from these correlations and midpoint BSM nodes in a quantum relay configuration have been predicted to offer higher tolerance to dark count effects [28].
+Given this, we predict that instead of Alice directly transmitting her qubit to Bob, our system can potentially offer higher tolerance to classical power in quantum-classical networks (see supp.
+materials).
+The coexistence study is conducted over spooled fiber in a laboratory located at Northwestern University’s Evanston campus.
 The transmission loss over the 30.2-km fiber is 4.9 dB at the classical signal’s 1547.32-nm wavelength and the 1290-nm quantum signals have a total loss of 10.1 dB (5.1 dB and 5.0 dB for Alice and Bob’s photons, respectively).
 At Charlie, the bypassing of the BSM by the classical signal using WDMs adds an additional 1.2 dB insertion loss.
-The total loss when transmitting the classical signal over the full 78.2-km fiber is 22.8 dB.
-Due to a receiver sensitivity of -18 dBm, the transceiver’s output power of -9 dBm requires amplification for the system to operate error-free.
-To achieve this, we place two cascaded erbium doped fiber amplifiers (EDFA) at the Evanston location just before multiplexing into the 30.2-km link to co-propagate with Alice’s photon.
-The minimum power launched into the 30.2-km fiber to achieve error-free communications could be as low as $-3$ dBm (0.5 mW) for the receiver located 30.2 km + 24 km away from the EDFA, and -10.7 dBm (85 $\mu$W) if the receiver were instead located directly at the end of the 30.2-km link.
-However, in this study we explore elevated launch powers to emulate more hostile environments that could be encountered in the field such as those found in high-capacity long-distance DWDM classical network systems.
+However, the classical transceiver is located at Northwestern’s Chicago campus which is connected to Evanston by a 24 km deployed fiber pair and thus operates over a total link distance of 30.2 + 48 km = 78.2 km.
+Due to the high loss over the full link (22.8 dB) and an initial output power of -9 dBm, amplification is needed to meet the classical system’s $-18$ dBm minimum received power requirements.
+We achieve this using erbium doped fiber amplifiers (EDFAs) placed just before multiplexing to co-propagate with Alice’s photon.
+The minimum power launched into the 30.2-km fiber to operate the 400-Gbps channel is $P_{\text{min}}=-3$ dBm (0.5 mW) for the receiver located 30.2 km + 24 km away from the EDFA, and -10.7 dBm (85 $\mu$W) if the classical system operated solely over the 30.2-km link.
+However, we explore elevated powers using the EDFAs to evaluate whether teleportation can coexist with more hostile classical systems such as ultra high-power long-distance DWDM conventional communications.
+Figure 2: (a) SpRS noise photon count rate in the Bell state measurement (BSM) detectors $D_{1}$ and $D_{2}$ as a function of the classical launch power $P_{\text{cl}}$ and the ratio $M$ to the minimum power $P_{\text{min}}$ needed to operate a single 400-Gbps channel.
+(b) Bob’s entanglement distribution visibility $V_{\rm ent}$ over 15.0 km to the BSM node versus $P_{\text{cl}}$ (top) and the corresponding two-fold coincidence count (CC) non-local interference fringe measured in the vertical (red) and anti-diagonal (blue) bases for $P_{\text{cl}}$ = 74 mW (bottom).
+(c) Hong-Ou-Mandel interference visibility $V_{\rm HOM}$ over 30.2 km between heralded photons from Alice and Bob’s sources versus $P_{\text{cl}}$ (top) and the corresponding four-fold coincidence interference fringe when $P_{\text{cl}}$ = 74 mW (bottom).
+The dashed lines show the predictions of our theoretical model [48] (see supp.
+materials).
 3 Experimental Results
@@
 2(a) shows the single photon count rates as we increase $P_{\text{cl}}$, where the SNSPD dark count rates have been subtracted.
-The extra noise scales linearly with power as 79.0 counts/s/mW and 97.9 counts/s/mW in $D_{1}$ and $D_{2}$, respectively, which differ due to slightly different detection system efficiencies.
-These results demonstrate that SpRS noise is significantly suppressed, as even mW power levels introduce extraneous noise on the order of the SNSPD dark count level.
-Since the 400-Gbps system can operate at $P_{\text{cl}}>0.5$ mW, this means that the SNSPDs hardly notice any rise in background counts when the system is operating at the minimum necessary classical power.
+The noise rates scale linearly as 79.0 counts/s/mW and 97.9 counts/s/mW in $D_{1}$ and $D_{2}$, respectively, the difference is due to slightly unequal efficiencies after the 50:50 coupler.
+Since a single 400-Gbps channel can operate with as low as $P_{\text{cl}}=0.5$ mW, the system can operate with nearly no noticeable noise increase above the SNSPD dark count rate.
+Fig.
+2(a) also includes an axis showing the ratio $M$ of $P_{\text{cl}}$ to the minimum single-channel power, which can be used to estimate the maximum aggregate data rates the $P_{\text{cl}}$ could support in a multi-channel system via $M\times 400$ Gpbs.
 As teleportation requires quantum entanglement as a resource and the ability to have indistinguishable photons for the BSM, we independently characterize the quality of Bob’s entanglement distribution to Charlie and HOM interference as we vary $P_{\text{cl}}$.
@@
 Fig.
-2(b) shows how Bob’s entanglement source’s non-local two-photon interference fringe visibility in the vertical and anti-diagonal bases is impacted after distributing one photon over 15.0 km to Charlie as we increase the classical power launched into the full 30.2-km link.
+2(b) shows how Bob’s entanglement source’s non-local two-photon interference fringe visibility in the vertical and anti-diagonal bases is impacted after distributing one photon over 15.0 km to Charlie as we increase the C-band signal’s power launched into the full 30.2-km link.
 The visibility is defined as $V_{\rm ent}=(R_{\rm max}-R_{\rm min})/(R_{\rm max}+R_{\rm min})$, where $R_{\rm max(\rm min)}$ are the maximum (minimum) two-fold coincidence counts of the interference fringe.
 Fig.
-2(b) shows no noticeable degradation of $V_{\rm ent}$ up to $P_{\text{cl}}=$ 74 mW, in which we measure visibilities of $V_{V}=97.5\pm 0.1\%$ and $V_{A}=95.3\pm 0.2\%$.
+2(b) shows no noticeable degradation of $V_{\rm ent}$ up to $P_{\text{cl}}=$ 74 mW, in which we measure $V_{\rm ent,V}=97.5\pm 0.1\%$ and $V_{\rm ent,A}=95.3\pm 0.2\%$.
 All error bars in this paper are determined via the Monte-Carlo method assuming Poisson photon counting statistics.
-These results indicate that entanglement distribution is well preserved at high classical powers, and that visibility is mainly limited from unity due to multi-pair emission during SPDC.
-These values are also well above the $1/\sqrt{2}$ bound for demonstrating the non-local nature of quantum entanglement [47].
-Thus, this state is more than sufficient to use as a resource for quantum teleportation.
+These results indicate that entanglement distribution is well preserved and that visibility is mainly limited by multi-pair emission during SPDC.
+These values are also well above the $1/\sqrt{2}$ bound for demonstrating the non-local nature of quantum entanglement [50] and more than sufficient to use for quantum teleportation.
 Next, we characterize HOM interference between heralded photons from Alice and Bob’s sources as we increase $P_{\text{cl}}$.
@@
 2(c).
-We measure a visibility of $V_{\rm HOM}=(R_{\rm max}-R_{\rm min})/R_{\rm max}=82.9\pm 4.5\%$ without the classical source, which is comparable to the $V_{\rm HOM}=83.1\pm 1.4\%$ visibility we obtain without distribution over the 30.2-km fiber.
+We measure a visibility of $V_{\rm HOM}=82.9\pm 4.5\%$ without the classical source, which is comparable to the $V_{\rm HOM}=83.1\pm 1.4\%$ visibility we obtain without distribution over the 30.2-km fiber (see supp.
+materials).
 We measure $V_{\rm HOM}=80.3\pm 3.8\%$ with $P_{\text{cl}}$ set to 74 mW, showing interference is not significantly degraded by high-power coexisting C-band light.
-Each visibility is well above the classical bound of 50% [48], indicating our photons are sufficiently indistinguishable for performing quantum teleportation.
-Figure 3: Four-fold coincidence counts as we scan Bob’s polarization analyzer setting along the Bloch sphere for Alice transmitting $|{D}\rangle$ (blue) and $|A\rangle$ (red) when the launch power of the 400-Gbps classical signal into the 30.2-km fiber is set to 74 mW.
-Figure 4: Real and imaginary components of the density matrices obtained by single-qubit quantum state tomography on Bob’s target photon, conditioned on a successful BSM, whilst the launch power of the 400-Gbps C-band signal into the 30.2-km fiber is set to 74 mW.
-(a-d) show the results for Alice transmitting $|\psi\rangle_{A}=|H\rangle$, $|V\rangle$, $|D\rangle$, and $|A\rangle$, respectively.
+Each visibility is well above the classical bound of 50% [51], indicating our ability to demonstrate non-classical interference across the range of the tested coexisting C-band powers.
 3.2 Quantum teleportation coexisting with high-rate classical communications
+Figure 3: Four-fold coincidence counts as we scan Bob’s polarization analyzer setting around the Bloch sphere for Alice transmitting $|{D}\rangle$ (blue) and $|A\rangle$ (red) when the launch power of the 400-Gbps C-band classical signal into the 30.2-km fiber is set to 74 mW.
 Looking at the performance of entanglement distribution and HOM interference, we find that the key operations underlying quantum teleportation are well maintained at high C-band power levels.
-To demonstrate that this translates to the ability to perform noise-resilient teleportation, we evaluate the teleportation of various qubits while simultaneously transmitting 74 mW of C-band classical power.
+To demonstrate that this translates to the ability to perform teleportation, we evaluate the teleportation of various qubits while simultaneously transmitting 74 mW of C-band classical power.
 Fig.
-3 shows four-fold coincidence counts as we scan Bob’s polarization analyzer setting along the Bloch sphere for Alice transmitting $|{D}\rangle=\frac{1}{\sqrt{2}}(|{H}\rangle+|{V}\rangle)$ and $|{A}\rangle=\frac{1}{\sqrt{2}}(|{H}\rangle-|{V}\rangle)$.
+3 shows four-fold coincidence counts as we scan Bob’s polarization analyzer setting around the Bloch sphere for Alice transmitting $|{D}\rangle=\frac{1}{\sqrt{2}}(|{H}\rangle+|{V}\rangle)$ and $|{A}\rangle=\frac{1}{\sqrt{2}}(|{H}\rangle-|{V}\rangle)$.
 The coincidences exhibit sinusoidal fringes that coincide with Alice’s prepared qubit.
-We obtain fringe visibilities of $V_{D}=81.3\pm 5.4\%$ and $V_{A}=74.7\pm 4.7\%$, both values being above the $V>1/3$ classical limit [49].
+We obtain fringe visibilities of $V_{D}=81.3\pm 5.4\%$ and $V_{A}=74.7\pm 4.7\%$, both values being above the $V=1/3$ classical limit [52].
 Next, we perform maximum likelihood quantum state tomography [53] on Bob’s target photon to reconstruct the density matrix $\rho_{B}$ of his quantum state conditioned on a BSM detection.
@@
 Knowing that SpRS noise rates are independent of Alice’s prepared qubit, assuming symmetry of the equatorial states we calculate the average fidelity for an arbitrary qubit via $F_{\rm avg}=(F_{H}+F_{V}+2(F_{D}+F_{A}))/6$ in which we obtain $F_{\rm avg}=89.9\pm 3.1\%$.
-Furthermore, we measure $F_{\rm avg}=90.4\pm 2.7\%$ without sending the photons over the 30.2-km link, demonstrating that there is minimal degradation of fidelity from distribution over the fiber link nor the inclusion of the classical light.
-All of the measured fidelities are shown in Fig.
-5.
-Since all are well above the classical bound of $F=2/3$ [49], we have clearly demonstrated that non-classical teleportation can be performed alongside high-rate conventional classical communications.
-Figure 5: Fidelity of the reconstructed density matrix of Bob’s state to Alice’s initially prepared state for four different qubits as well as the average fidelity when the power of the 400-Gbps classical signal is set to 74 mW.
-The horizontal line shows the $F>2/3$ limit for demonstrating non-classical teleportation.
+Furthermore, we measure $F_{\rm avg}=90.8\pm 0.8\%$ without sending the photons over the 30.2-km link (see supp.
+materials), demonstrating that there is minimal degradation of fidelity from distribution over the fiber nor the inclusion of the 400-Gbps signal.
+The state transfer fidelity of each qubit is shown in Fig.
+4(e).
+All are well above the $F=2/3$ limit for classical-based methods [52], demonstrating that non-classical teleportation can be achieved alongside high-rate conventional communications.
+Figure 4: Real and imaginary components of the density matrices obtained by single-qubit quantum state tomography on Bob’s target photon, conditioned on a successful BSM, whilst the launch power of the 400-Gbps C-band signal into the 30.2-km fiber is set to 74 mW.
+(a-d) show the results for Alice transmitting $|\psi\rangle_{A}=|H\rangle$, $|V\rangle$, $|D\rangle$, and $|A\rangle$, respectively.
+(e) The fidelity of Bob’s density matrix to Alice’s ideally prepared qubit and the calculated average fidelity.
+The dotted line shows the $F=2/3$ limit for the highest fidelity achievable using only classical physics-based methods [52].
 4 Discussion and Outlook
-Since teleportation was conducted alongside much higher powers than necessary for the 400-Gbps classical system, this indicates that substantially higher data rates could have propagated through the fiber.
-The classical power used was 148 times greater than necessary for a single 400-Gbps channel, and 870 times greater than if the receiver were located directly at the end of the 30.2-km fiber.
-Thus, a multi-channel WDM classical system could easily have carried many terabits/s aggregate data rates without having a significant impact on quantum fidelity.
-The demonstrated powers are also on the order of the highest achieved in state-of-the-art single [10, 11, 12] and entangled [17] photon coexistence studies.
-This indicates the potential for teleportation-based quantum networking systems, such as quantum repeaters, operating alongside classical backbone networks [10].
-Since our system uses two different photon pair sources, multiple nodes, and four-fold coincidence detection, our results indicate the potential for more complex systems operating alongside classical networks.
-For example, simple modifications to Alice’s source could allow entangled photon pair generation.
-A BSM at Charlie would then result in entanglement swapping [52], establishing entanglement between initially independent photons at Alice and Bob’s nodes.
-Due to our system’s high tolerance to classical power, it is reasonable to conclude that a similarly designed entanglement swapping system could operate alongside high-rate coexisting classical communications.
-Further, our multi-node configuration with the BSM performed near the midpoint of a longer fiber link is a key element for multi-node networking with a central quantum processor [34] and to implement quantum repeaters [29].
-We note that these applications will also require quantum memories to store incident photons, the degree to which SpRS noise will impact various platforms and applications will require future investigation.
-The state teleportation fidelity we obtained is comparable to recent experiments conducted over quantum-dedicated fiber networks [34, 33, 37], but could be further improved via standard methods.
-We showed that opposed to SpRS noise from coexisting classical light, our fidelity was mainly limited by multi-photon pair emission during SPDC and imperfect photon indistinguishably during the BSM.
-Raw fidelity could thus be improved by reducing the photon pair generation probability [46] and increasing indistinguishability by decreasing the ratio between the source’s pump spectral bandwidth to the photon filter bandwidths [42].
-When considering SpRS noise, narrower filters would simultaneously increase indistinguishability and reject more SpRS noise.
-Altogether, employing these approaches could allow our teleportation fidelity to approach more closely to unity and have an even higher tolerance to classical power.
-Finally, classical signals are often used for quantum networking purposes.
-Some applications include communicating measurement results between nodes (eg., the BSM), synchronizing distant pulsed sources and detectors to a common clock, and channel monitoring.
-Similarly, quantum-classical coexistence can enhance functionality without sacrificing potentially limited fiber resources [16, 53, 22, 21].
-Since these signals can typically operate well below the powers demonstrated here (sub-mW levels), our results show that including these signals in our system would have negligible impact if allocated to the C- or L-bands.
-In conclusion, we have demonstrated a three-node quantum state teleportation system operating in a 30.2-km fiber link that simultaneously carries 400-Gbps conventional classical communications.
-By employing various methods to suppress the impact of SpRS noise photons, teleportation fidelity was well maintained with elevated classical powers capable of transmitting many terabits/s aggregate data rates.
-Since we have clearly shown that appropriately designed teleportation systems can operate alongside modern high-rate data traffic, this work demonstrates a significant step towards ensuring future quantum networking can be realized anywhere in the existing fiber infrastructure.
+In this paper, teleportation was performed alongside $M\approx$150 times higher classical powers than necessary for error-free 400-Gbps communications over the 78-km fiber and $M\approx$870 times greater for a 30.2-km link.
+Thus, WDM communications could have easily reached many Tbps aggregate data rates without significant impact on teleportation fidelity.
+The classical power was also on the order of the highest achieved in state-of-the-art studies using direct transmission of quantum states [25, 10, 11, 12, 17] and indicates the potential for teleportation-based applications operating in the classical backbone fiber infrastructure [10].
+Our results also indicate the potential for integrating even more advanced quantum operations.
+For example, simple modifications to Alice’s photon pair source would generate entangled photon pairs and result in entanglement swapping [55], entangling initially independent photons at Alice and Bob’s nodes after the BSM.
+Based on this similarity, it is reasonable to conclude that swapping alongside high-rate classical data is achievable.
+Further, midpoint BSMs are important for applications such as networking with a central quantum processor [34], quantum repeaters [29], amongst others.
+Although these will require more capabilities beyond those demonstrated here (eg., quantum memories, purification, etc.), our results are promising and could be useful for the ongoing development of this technology.
+Although O-band teleportation networks show promise for high-fidelity integration, C-band quantum networking is usually preferred in quantum-dedicated fiber due to the lowest loss in silica fibers.
+Unfortunately, C-band channels suffer from orders of magnitude higher SpRS generation rates from C-band classical light [4, 17].
+CV quantum encoding can help address this issue [25] but is much more fragile to loss compared to discrete-variables which limits its reach.
+One intriguing possibility is to use high-dimensional entanglement [56], but has yet to be fully investigated in high-power regimes.
+Using a theoretical model [48], we predict our teleportation fidelity would be significantly impacted with mW-level C-band classical power if we used C-band quantum signals (see supp.
+materials), but coincidence rates would roughly triple.
+This has relevance for research towards long-distance quantum backbone networks since conventional backbone fibers can have $\sim 100$ mW aggregate power.
+Further, WDM bandwidth may become limited as some conventional networks push towards fully populating the C- and L-bands [57], which also increases the difficulty of avoiding channel cross-talk and optical amplifier noise.
+One route to alleviate lower O-band quantum rates is to use wavelength multiplexing, which has been shown successful in allowing multi-channel entangled photons populating most of the O-band co-propagating with high-power C- and L-band WDM channels [17].
+In any case, future research into novel methods to protect quantum fidelity from strong background noise would be invaluable.
+Some teleportation-based applications require storing photons in quantum memories (QMs), for example to store target photons long enough to receive a signal encoding the BSM result and apply the corresponding unitary transformation to recover Alice’s exact state before usage.
+In our experiment we performed a posteriori teleportation [58] since Bob’s photon is detected before the BSM occurs.
+Interestingly, quantum mechanics predicts the same correlations irrespective of measurement order.
+However, attempting to store photons over a long duration can increase photon loss which may complicate the dependence on measurement order in real-world settings having appreciable background noise.
+If Bob’s target channel is completely noiseless, our model predicts that adding pure loss to his target photon’s channel does not further degrade the fidelity set by the noise level in the BSM detectors.
+However, this no longer holds if there is some noise count probability in Bob’s detector.
+In our case we predict our low dark count SNSPDs could tolerate roughly 60 dB loss in the target channel, which is more than manageable using spooled fiber to delay measurement.
+However, other approaches such as those using quantum frequency conversion to interface telecom wavelength photons with particular QMs can have notable noise-levels [37] that place more stringent requirements on Bob’s channel efficiency (see supp.
+materials).
+It is interesting to consider whether teleporting quantum information has any unique benefits in quantum-classical networking compared to direct transmission.
+Interestingly, we predict that our system can achieve higher coexisting classical power compared to if Alice had directly sent her qubit to Bob over the fiber using a WCS or a heralded single-photon (see supp.
+materials).
+This could be due to the stronger correlations in detection across the four photons in our system, and also the segmenting of the longer fiber with a midpoint BSM [28].
+Future studies will compare with other teleportation designs such as those using a WCS source at Alice’s node [33, 36, 39] instead of our heralding approach.
+Looking beyond state teleportation, it would be interesting to investigate the impact of coexisting classical channels on concatenated entanglement swapping or quantum repeaters.
+In regards to real-world deployment, classical light-based signals are often used in quantum networking for synchronizing distant pulsed quantum sources and detectors to a common clock, communicating measurement results (eg., BSM results), and monitoring environmental disturbances.
+Similarly, multiplexing these signals can enhance functionality without sacrificing potentially limited fiber resources [16, 59, 22, 21].
+Since these can typically operate at sub-mW powers, our results show that including these in our system would have negligible impact if allocated to the C- or L-bands.
+In conclusion, we have demonstrated quantum state teleportation over a 30.2-km fiber that is populated with high-power 400-Gbps conventional data traffic.
+By employing various methods to suppress SpRS noise, teleportation fidelity was well maintained alongside elevated classical powers capable of transmitting many Tbps aggregate data rates.
+We further investigated multiple key questions for optimizing teleportation-based applications and identified challenges facing the deployment of future quantum networks.
+Altogether, this work demonstrates a significant step towards ensuring complex multi-photon/multi-node quantum network applications can be realized anywhere in the existing fiber infrastructure.
 […]
@@
 Acknowledgments The authors would like to acknowledge the support of StarLight International/National Communications Exchange Facility and Ciena Corporation.
-We further thank the members of the Advanced Quantum Networks for Scientific Discovery (AQNET-SD) project, which is a collaboration between Northwestern University, Caltech, Fermilab National Laboratory, Argonne National Laboratory, the Jet Propulsion Laboratory, and University of Illinois Urbana-Champaign towards advancing quantum network technology and applications.
+We further thank the members of the Advanced Quantum Networks for Scientific Discovery (AQNET-SD) project, which is a collaboration between Northwestern University, Caltech, Fermilab, Argonne National Laboratory, Jet Propulsion Laboratory, and University of Illinois Urbana-Champaign.
+We also thank Gina Talcott for help in the editing process.
 […]

### v4

@@
 To protect quantum fidelity from spontaneous Raman scattering noise, we use optimal O-band quantum channels, narrow spectro-temporal filtering, and multi-photon coincidence detection.
-Fidelity is shown to be well maintained with an elevated C-band classical power of 18.7 dBm, which could support multiple classical channels totaling many terabits/s aggregate data rates.
+Fidelity is shown to be well maintained with an elevated C-band launch power of 18.7 dBm for the single-channel 400-Gbps signal, which we project could support multiple classical channels totaling many terabits/s aggregate data rates.
 These results show the feasibility of advanced quantum and classical network applications operating within a unified fiber infrastructure.
@@
 Without careful design, a trade-off between the capabilities of conventional and quantum networking arises.
-The study of quantum-classical coexistence has a long history with many experiments conducted for applications using weak coherent state (WCS) sources [3, 4, 5, 6, 7, 8, 9, 10, 11, 12], entangled photon pairs [13, 14, 15, 16, 17, 18, 19, 20, 21, 22], Bell state measurements (BSMs) on WCSs [23, 24], continuous variables [25], and squeezed light [26].
+The study of quantum-classical coexistence has a long history with many experiments conducted for applications using weak coherent state (WCS) sources [3, 4, 5, 6, 7, 8, 9, 10, 11, 12], entangled photon pairs [13, 14, 15, 16, 17, 18, 19, 20, 21, 22], Bell state measurements (BSMs) on WCSs [23, 24], continuous variables (CVs) [25], and squeezed light [26].
 Although significant achievements and insights have been made in these studies, all experiments to date have focused on systems that directly transmit quantum information between network nodes.
@@
 Teleportation is performed via a joint BSM on a single photon and one member of an entangled Bell state photon pair [40].
-Utilizing wavelength multiplexing of quantum and classical signals, a heralded single photon carrying a polarization encoded qubit and one photon of an entangled Bell state photon pair are distributed over roughly 15 km of fiber to undergo a BSM near the midpoint of the fiber, whilst classical light traverses the full 30.2 km link plus an additional 48 km of deployed fiber.
+Utilizing wavelength multiplexing of quantum and classical signals, a heralded single photon carrying a polarization encoded qubit and one photon of an entangled Bell state photon pair are distributed over roughly 15 km of spooled fiber to undergo a BSM near the midpoint of the fiber, whilst classical light traverses the full 30.2-km link plus an additional 48 km of deployed fiber.
 We implement multiple techniques to mitigate degradation due to noise photons.
@@
 Furthermore, each source uses spontaneous parametric down-conversion to generate photon pairs with strong temporal correlations to further reject noise via coincidence detection of the four photons.
-The design is also similar to a quantum relay, which has been predicted to reduce the impact of detector dark count noise [28] indicating a potential application in quantum-classical coexistence.
 We first demonstrate that our approach allows high-power C-band communications to have no significant impact on entanglement distribution to the BSM node nor Hong-Ou-Mandel (HOM) interference [41] between independent heralded photons, which are the underlying operations of the quantum teleportation protocol [40].
 We then perform quantum state teleportation and demonstrate that fidelity is maintained at elevated C-band launch powers as high as 74 mW (18.7 dBm), which is on the order of the highest achieved using direct transmission of quantum states [10, 11, 12, 17, 25].
-Using a theoretical model, we then discuss the the pros and cons of using noisier quantum wavelength channels (eg., C-band), the effects of storage in imperfect quantum memories, and compare to methods using direct qubit transmission.
+Using a theoretical model, we then discuss the the pros and cons of using noisier quantum wavelength channels (e.g., C-band), the effects of storage in imperfect quantum memories, and compare to methods using direct qubit transmission.
 This experiment demonstrates approaches for integrating many key classical and quantum network components, including high-rate classical signals, optical amplifiers, single and entangled photon pair distribution, and multi-qubit operations such as BSMs all operating simultaneously in the same fibers.
@@
 Two photons generated at Bob’s node are prepared in an entangled Bell state, where one is transmitted to counter-propagate with the classical light to Charlie to undergo a Bell state measurement (BSM) with Alice’s qubit.
-After a BSM detection, Alice’s state is destroyed whilst the photon kept at Bob’s node is projected onto the state $\sigma|\psi\rangle_{A}$, where $\sigma$ is a unitary Pauli transformation unique to the BSM result.
+After a BSM detection, Alice’s state is destroyed whilst the photon kept at Bob’s node is projected onto the state $\sigma|\psi\rangle_{A}$, where $\sigma$ is a unitary Pauli transformation identifiable from the BSM result.
 (b) Experimental implementation.
-Quantum and classical signals coexist in 30.2  km of fiber ($L_{AC}=15.2\text{ km}$, $L_{BC}=15.0\text{ km}$).
+Quantum and classical signals coexist in 30.2  km of spooled fiber ($L_{AC}=15.2\text{ km}$, $L_{BC}=15.0\text{ km}$).
 Alice and Bob both generate photon pairs via non-degenerate spontaneous parametric down-conversion with center wavelengths of 1290 nm and 1310 nm.
 Alice’s source heralds a single photon, whereas Bob prepares a two-photon entangled Bell state.
-Qubits are encoded in polarization and the 1290-nm photons from each source are transmitted to Charlie for the BSM and teleportation is evaluated via four-fold coincidence detection and polarization analysis at Bob.
-The classical signal has a rate of 400-Gbps and is allocated to the 1547.32 nm C-band wavelength channel.
-This signal travels an additional 24-km of deployed fiber before and after multiplexing giving a total classical link distance of 78.2 km (FPC = fiber polarization controller, DWDM = dense wavelength division multiplexer, FWDM = O-band/C-band WDM, PPLN = periodically poled lithium niobate waveguide, $\lambda/2$ = half-wave plate, $\lambda/4$ = quarter-wave plate, LP = linear polarizer, LCR = liquid crystal retarder, PBS = polarizing beam splitter, FBS = 50:50 fiber coupler, VDL = variable optical delay line, FBG = fiber Bragg grating, CIR = circulator, $\text{D}_{j}$ = superconducting nanowire single photon detector, C = common port, P = pass port, R = reflect port, EDFA = erbium doped fiber amplifier).
+Qubits are encoded in polarization and the 1290-nm photons from each source are transmitted to Charlie for the BSM.
+Teleportation is evaluated via four-fold coincidence detection and polarization analysis of Bob’s target photon.
+The classical source transmits a 400-Gbps C-band signal (1547.32 nm) over 24 km of deployed fiber before being amplified and multiplexed into the 30.2-km spool.
+After de-multiplexing this propagates over another 24-km deployed fiber to the receiver, totaling 78.2 km of fiber.
+(FPC = fiber polarization controller, DWDM = dense wavelength division multiplexer, FWDM = O-band/C-band WDM, PPLN = periodically poled lithium niobate waveguide, $\lambda/2$ = half-wave plate, $\lambda/4$ = quarter-wave plate, LP = linear polarizer, LCR = liquid crystal retarder, PBS = polarizing beam splitter, FBS = 50:50 fiber coupler, VDL = variable optical delay line, FBG = fiber Bragg grating, CIR = circulator, $\text{D}_{j}$ = superconducting nanowire single photon detector, C = common port, P = pass port, R = reflect port, EDFA = erbium doped fiber amplifier).
 2 Experimental design
@@
 materials).
-The 650-nm pulses generated inside each waveguide simultaneously drives SPDC to produce a two-mode squeezed vacuum state with a non-degenerate joint spectrum centered around 1300 nm.
-This state approximates a probabilistic source of single-photon pairs given the probability of generating a pair per pulse is $\ll 1$.
+The 650-nm pulses generated inside each waveguide simultaneously drives SPDC to produce a superposition of photon pairs with a non-degenerate spectrum centered around 1300 nm and forms the basis for Alice’s heralded single-photon source (HSPS) and generating Bob’s entangled photons, but requires a low single-photon pair generation rate to minimize the probability of multi-photon emission [43]
 Alice’s qubit is initially encoded onto a heralded single photon.
@@
 Including the PBSs thus allows us to investigate BSM nodes most likely to be implemented in future networks.
-We also note that due to polarization mode dispersion over long fibers, the SpRS noise is largely unpolarized [46, 17] which means the polarizing elements roughly halve the amount of SpRS entering each detector.
+We note that SpRS photons in long-distance fibers are often unpolarized due to polarization mode dispersion [47, 17] meaning polarizing elements roughly halve the noise rates.
+This was confirmed in our system by polarization analysis of the noise photons at the BSM node (see supp.
+materials).
 […]
 Since the classical data modulation is uncorrelated with the pulsed quantum sources in the time-domain, the coincidence detection window acts as a temporal filter of the SpRS noise.
-The temporal correlations in photon pair generation can also be exploited for time-domain filtering.
-Heralding single photons can improve tolerance to noise compared to sources such as WCSs by accepting a detection only when the heralding signal ensures a photon had been generated which reduces the probability of transmitting a vacuum state [47].
-In our four-photon teleportation system, both Alice and Bob’s pair sources benefit from these correlations and midpoint BSM nodes in a quantum relay configuration have been predicted to offer higher tolerance to dark count effects [28].
-Given this, we predict that instead of Alice directly transmitting her qubit to Bob, our system can potentially offer higher tolerance to classical power in quantum-classical networks (see supp.
-materials).
+The temporal and photon-number correlations of our photon-pair sources further help the system’s tolerance to noise by accepting detection events at the receiver only in the time slots when a detection in the heralding arm indicated a photon was transmitted over the noisy channel [48].
+This can significantly improve the signal-to-noise ratio compared to non-heralded sources that have a high probability of transmitting a vacuum state such as WCSs.
+In our four-photon teleportation system, both Alice and Bob’s pair sources benefit from these correlations.
 The coexistence study is conducted over spooled fiber in a laboratory located at Northwestern University’s Evanston campus.
@@
 At Charlie, the bypassing of the BSM by the classical signal using WDMs adds an additional 1.2 dB insertion loss.
-However, the classical transceiver is located at Northwestern’s Chicago campus which is connected to Evanston by a 24 km deployed fiber pair and thus operates over a total link distance of 30.2 + 48 km = 78.2 km.
+However, the classical transceiver is located at Northwestern’s Chicago campus which is connected to Evanston by a 24-km deployed fiber pair and thus operates over a total link distance of 30.2 + 48 km = 78.2 km.
 Due to the high loss over the full link (22.8 dB) and an initial output power of -9 dBm, amplification is needed to meet the classical system’s $-18$ dBm minimum received power requirements.
@@
 Fig.
-2(a) also includes an axis showing the ratio $M$ of $P_{\text{cl}}$ to the minimum single-channel power, which can be used to estimate the maximum aggregate data rates the $P_{\text{cl}}$ could support in a multi-channel system via $M\times 400$ Gpbs.
+2(a) also shows the ratio $M=P_{\text{cl}}/P_{\text{min}}$, which is often used to roughly estimate the maximum number of channels $M$ that $P_{\text{cl}}$ could support across multiple WDM channels (e.g., [23, 24]) assuming each operates at $P_{\text{min}}$ and has similar SpRS properties (see supp.
+materials).
 As teleportation requires quantum entanglement as a resource and the ability to have indistinguishable photons for the BSM, we independently characterize the quality of Bob’s entanglement distribution to Charlie and HOM interference as we vary $P_{\text{cl}}$.
@@
 2(b) shows no noticeable degradation of $V_{\rm ent}$ up to $P_{\text{cl}}=$ 74 mW, in which we measure $V_{\rm ent,V}=97.5\pm 0.1\%$ and $V_{\rm ent,A}=95.3\pm 0.2\%$.
-All error bars in this paper are determined via the Monte-Carlo method assuming Poisson photon counting statistics.
+All values are determined by raw data without accidental count subtraction and error bars are calculated via the Monte-Carlo method assuming Poisson photon counting statistics.
 These results indicate that entanglement distribution is well preserved and that visibility is mainly limited by multi-pair emission during SPDC.
@@
 materials).
-We measure $V_{\rm HOM}=80.3\pm 3.8\%$ with $P_{\text{cl}}$ set to 74 mW, showing interference is not significantly degraded by high-power coexisting C-band light.
+We measure $V_{\rm HOM}=80.3\pm 3.8\%$ with $P_{\text{cl}}$ set to 74 mW, which shows that interference is well preserved in the presence of high-power C-band light but is still limited by the imperfect spectral and single-photon purity of our sources.
 Each visibility is well above the classical bound of 50% [51], indicating our ability to demonstrate non-classical interference across the range of the tested coexisting C-band powers.
@@
 3 shows four-fold coincidence counts as we scan Bob’s polarization analyzer setting around the Bloch sphere for Alice transmitting $|{D}\rangle=\frac{1}{\sqrt{2}}(|{H}\rangle+|{V}\rangle)$ and $|{A}\rangle=\frac{1}{\sqrt{2}}(|{H}\rangle-|{V}\rangle)$.
-The coincidences exhibit sinusoidal fringes that coincide with Alice’s prepared qubit.
+We observe for each case a sinusoidal fringe that resembles Alice’s prepared qubit with a maximum four-fold rate of $\sim$0.09 counts per second.
 We obtain fringe visibilities of $V_{D}=81.3\pm 5.4\%$ and $V_{A}=74.7\pm 4.7\%$, both values being above the $V=1/3$ classical limit [52].
@@
 The fidelity of $\rho_{B}$ to Alice’s ideal qubit $|\psi\rangle_{A}$ is determined by calculating $F_{\psi}=\langle\psi_{A}|\rho_{B}|\psi_{A}\rangle$.
-State transfer in the H/V basis does not require HOM interference due to the projection onto $|H\rangle_{D_{1}}|V\rangle_{D_{2}}$ in the Bell state analyzer [54], where raw fidelity is mainly limited from unity due to multi-photon pair degradation.
+State transfer in the $H/V$ basis does not require HOM interference due to the projection onto the qubit eigen-states in the Bell state analyzer [54], where raw fidelity is mainly limited from unity due to multi-photon pair degradation.
 In this basis, we measure $F_{H}=97.5\pm 1.2\%$ and $F_{V}=95.8\pm 2.5\%$ for $|{\psi}\rangle_{A}=|{H}\rangle$ and $|{V}\rangle$, respectively.
@@
 For Alice transmitting $|{\psi}\rangle_{A}=|{D}\rangle$ and $|{A}\rangle$, we obtain fidelities of $F_{D}=87.5\pm 3.9\%$ and $F_{A}=85.5\pm 3.7\%$, respectively.
-Knowing that SpRS noise rates are independent of Alice’s prepared qubit, assuming symmetry of the equatorial states we calculate the average fidelity for an arbitrary qubit via $F_{\rm avg}=(F_{H}+F_{V}+2(F_{D}+F_{A}))/6$ in which we obtain $F_{\rm avg}=89.9\pm 3.1\%$.
+Since the SpRS noise is unpolarized and independent of Alice’s prepared qubit, assuming symmetry of the remaining equatorial states we obtain an average fidelity for an arbitrary qubit on the Bloch sphere of $F_{\rm avg}=\frac{1}{3}F_{\text{poles}}+\frac{2}{3}F_{\text{equator }}=89.9\pm 3.1\%$.
 Furthermore, we measure $F_{\rm avg}=90.8\pm 0.8\%$ without sending the photons over the 30.2-km link (see supp.
@@
 Further, midpoint BSMs are important for applications such as networking with a central quantum processor [34], quantum repeaters [29], amongst others.
-Although these will require more capabilities beyond those demonstrated here (eg., quantum memories, purification, etc.), our results are promising and could be useful for the ongoing development of this technology.
+Although these will require more capabilities beyond those demonstrated here (e.g., quantum memories, purification, etc.), our results are promising and could be useful for the ongoing development of this technology.
 Although O-band teleportation networks show promise for high-fidelity integration, C-band quantum networking is usually preferred in quantum-dedicated fiber due to the lowest loss in silica fibers.
@@
 One intriguing possibility is to use high-dimensional entanglement [56], but has yet to be fully investigated in high-power regimes.
-Using a theoretical model [48], we predict our teleportation fidelity would be significantly impacted with mW-level C-band classical power if we used C-band quantum signals (see supp.
-materials), but coincidence rates would roughly triple.
+Using our theoretical model [49], we predict C-band teleportation fidelity would be significantly degraded with mW-level coexisting C-band power (see supp.
+materials) but would improve our fiber transmission loss by $\sim 0.15$ dB/km.
 This has relevance for research towards long-distance quantum backbone networks since conventional backbone fibers can have $\sim 100$ mW aggregate power.
@@
 It is interesting to consider whether teleporting quantum information has any unique benefits in quantum-classical networking compared to direct transmission.
-Interestingly, we predict that our system can achieve higher coexisting classical power compared to if Alice had directly sent her qubit to Bob over the fiber using a WCS or a heralded single-photon (see supp.
+It has been shown that using teleportation and BSM operations periodically along a long-distance fiber can increase the possible distances for quantum applications limited by detector dark count noise [28], however this may not necessarily apply for SpRS noise generated by classical light and depends on a vast number of parameters such as wavelength channel selections, fiber lengths, co- or counter-propagating signals, and optical amplification for long-distance classical communications.
+We predict that in some cases there is no advantage if the system has high SpRS noise, but using the methods applied here can help recover some improvements over long-distances and could in principle operate over 100’s of kms, albeit at impractically low quantum rates (see supp.
 materials).
-This could be due to the stronger correlations in detection across the four photons in our system, and also the segmenting of the longer fiber with a midpoint BSM [28].
-Future studies will compare with other teleportation designs such as those using a WCS source at Alice’s node [33, 36, 39] instead of our heralding approach.
-Looking beyond state teleportation, it would be interesting to investigate the impact of coexisting classical channels on concatenated entanglement swapping or quantum repeaters.
-In regards to real-world deployment, classical light-based signals are often used in quantum networking for synchronizing distant pulsed quantum sources and detectors to a common clock, communicating measurement results (eg., BSM results), and monitoring environmental disturbances.
+A full analysis of the various scenarios to consider will be pursued in future work.
+We will also investigate teleportation using a WCS source at Alice’s node [33, 36, 38, 39] as well as the consequences of multiplexing classical communications with entanglement swapping, quantum repeaters, and multi-photon entanglement.
+In regards to real-world deployment, classical light-based signals are often used in quantum networking for synchronizing distant pulsed quantum sources and detectors to a common clock, communicating measurement results (e.g., BSM results), and monitoring environmental disturbances.
 Similarly, multiplexing these signals can enhance functionality without sacrificing potentially limited fiber resources [16, 59, 22, 21].